OpenStudy (anonymous):

Help! How would I solve tan60=6/x

OpenStudy (anonymous):

tan (60)=6/x x=6/tan(60)

OpenStudy (anonymous):

Would i do 6 divided by 3 and then divide by the 60? @AJ01

OpenStudy (anonymous):

no x=6/(tan(60))

OpenStudy (anonymous):

tan is a function take tan (60)

OpenStudy (anonymous):

Would I do 3 * 60 then divide by 6?

OpenStudy (anonymous):

Would I have to plug in a number for tan?

OpenStudy (anonymous):

yes 6/tan (60)=6/1.732

OpenStudy (anonymous):

Would I do 60*1.732? Then divide?

OpenStudy (anonymous):

Or 6/1.732 then *60

OpenStudy (anonymous):

@AJ

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

x=6/1.732=3.464

OpenStudy (anonymous):

Then would I take the 3.464 and *60?

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

Or would I divide the 60 and 3.464 and get 17?

OpenStudy (anonymous):

@AJ01